### 04 Mar 2026 v1

# Tight inapproximability of max-LINSAT and implications for decoded quantum interferometry

Abstract
We establish tight inapproximability bounds for max-LINSAT, the problem of maximizing the number of satisfied linear constraints over the finite field $\mathbb{F}_{q}$, where each constraint accepts $r$ values. Specifically, we prove by a direct reduction from Håstad’s theorem that no polynomial-time algorithm can exceed the random-assignment ratio $r/q$ by any constant, assuming $\mathsf{P}\neq\mathsf{NP}$. This threshold coincides with the $\ell/m\to 0$ limit of the semicircle law governing decoded quantum interferometry (DQI), where $\ell$ is the decoding radius of the underlying code: as the decodable structure vanishes, DQI’s approximation ratio degrades to exactly the worst-case bound established by our result. Together, these observations delineate the boundary between worst-case hardness and potential quantum advantage, showing that any algorithm surpassing $r/q$ must exploit algebraic structure specific to the instance.
1 Introduction
Recent years have seen growing interest in quantum algorithms for combinatorial optimization [1], particularly for problems with strong worst-case classical hardness. Given the practical and industrial relevance of these problems across a wide range of scheduling and routing tasks, the design of corresponding quantum algorithms is of considerable importance. For such problems, one cannot expect a quantum computer to solve or approximate all instances of the combinatorial optimization problem in polynomial time. This does not mean that one cannot hope that quantum computers will have computational advantages over classical computers. Indeed, in such cases, potential super-polynomial quantum advantages are often investigated on restricted or structured families of instances [33, 42, 41, 12]. One such example that has recently been given a lot of attention is the decoded quantum interferometry (DQI) algorithm introduced in Ref. [25]. DQI works by using the quantum Fourier transform to reduce the optimization problem to a decoding task, and then coherently applies an efficient decoding algorithm to amplify the quantum interference between satisfying assignments. The approximation ratio of DQI is characterized by a closed-form formula, coined the semicircle law [25].
DQI targets a generalized linear constraint satisfaction problem over finite fields, commonly referred to as max-LINSAT. The flagship application is optimal polynomial intersection (OPI) problem, a structured max-LINSAT instance whose constraint matrix has Vandermonde structure arising from Reed–Solomon codes that are efficiently decodable up to their designed error radius. For OPI, in a certain parameter regime, DQI achieves a super-polynomial speed-up over known classical algorithms.
A growing body of follow-up work has developed around DQI [14, 32, 37, 10, 9, 39, 30, 2, 34, 31, 19, 28, 40, 23, 26, 15, 38, 11]. Most relevant to our setting, the DQI advantage for OPI has been strengthened via soft decoding of Reed–Solomon codes [14, 15] and extended to algebraic geometry codes via a generalization called HOPI [19]. At the same time, several works have exposed DQI’s limitations [31, 2, 30].
A prerequisite for precisely interpreting the results surrounding DQI is a clear understanding of the approximability limits of the underlying optimization problem. For many well-studied constraint satisfaction problems, tight inapproximability thresholds are known, often matching the performance of a random assignment. For max-LINSAT in the formulation considered by DQI, which allows linear constraints with arbitrary acceptance sets over a finite field, no such worst-case guarantees have previously been established. It is, therefore, unclear whether the approximation ratios achieved by quantum or classical heuristics on generic instances are close to being optimal, or whether substantially better algorithms might exist.
We address this gap in the present work by establishing tight inapproximability results for max-LINSAT over finite fields. We consider max-LINSAT instances over a finite field [^1] Recall that the finite field $\mathbb{F}_{q}$ of order $q$ exists whenever $q=p^{\tau}$ for some prime $p$ and integer $\tau\geq 1$. $\mathbb{F}_{q}$ in which each constraint is satisfied whenever a linear form evaluates to one of $r$ allowed values, with $1\leq r\leq q-1$. This class includes, as a special case, the linear constraint satisfaction problems studied in Ref. [25]. A random assignment satisfies each constraint with probability exactly $r/q$, yielding a natural baseline approximation ratio.
Our main result shows that this random baseline is provably optimal in the worst case. Specifically, for every finite field $\mathbb{F}_{q}$ of order $q$, every $1\leq r\leq q-1$, and every $\varepsilon>0$, it is $\mathsf{NP}$-hard to distinguish between instances in which almost all constraints are satisfiable and instances in which no assignment satisfies more than a $(r/q+\varepsilon)$-fraction of constraints. Assuming $\mathsf{P}\neq\mathsf{NP}$, no polynomial-time classical algorithm can approximate max-LINSAT beyond the random-assignment threshold, and likewise no polynomial-time quantum algorithm can do so unless $\mathsf{NP}\subseteq\mathsf{BQP}$. From the perspective of quantum optimization, this provides a complexity-theoretic benchmark. The threshold $r/q$ is hence a hard wall for worst-case max-LINSAT(q,r)(q,r): no efficient algorithm can consistently exceed it on arbitrary instances. This does not conflict with DQI’s demonstrated advantage on OPI, where the Vandermonde structure of classical Reed–Solomon codes enables approximation ratios far above the random baseline. Rather, our result delineates the boundary: any algorithm that beats $r/q$ on a max-LINSAT instance must exploit structure specific to that instance. The potential for quantum advantage in this setting lies in leveraging algebraic structure rather than overcoming worst-case complexity barriers.
The proof builds on and generalizes an inapproximability result of Håstad for max-E33-LIN over finite Abelian groups, which establishes tight hardness for the case of single-valued acceptance sets [21]. We show how this hardness extends to max-LINSAT instances with uniform acceptance sets of size $r$, yielding a matching hardness threshold for the general problem. The resulting inapproximability bounds are tight.
2 Preliminaries
We start by defining the optimization problems considered in this work, in particular focusing on the max-LINSAT-problem as defined in Ref. [25].
Definition 1 (max-LINSAT over $\mathbb{F}_{q}$).
Let $\mathbb{F}_{q}$ be the finite field of order $q$ and let $B\in\mathbb{F}_{q}^{m\times n}$. For each $i=1,\ldots,m$, let $F_{i}\subset\mathbb{F}_{q}$ be an arbitrary subset of $\mathbb{F}_{q}$, which yields a corresponding constraint $\sum_{j=1}^{n}B_{i,j}x_{j}\in F_{i}$. The max-LINSAT task is to find $\boldsymbol{x}\in\mathbb{F}_{q}^{n}$ satisfying as many as possible of these $m$ constraints. We denote by $\mathrm{OPT}$ the maximum number of satisfiable constraints and define an approximation ratio $\alpha$ for a solution satisfying $s$ out of $m$ constraints to be $\alpha\coloneq s/m$.
We assume that the acceptance sets $F_{i}$ are given as an explicit list.
Definition 2 (max-LINSAT(q,r)(q,r)).
We write max-LINSAT(q,r)(q,r) for the restriction of max-LINSAT over $\mathbb{F}_{q}$ to instances in which $|F_{i}|=r$ for every $i$. The special case max-XORSAT is max-LINSAT(2,1)(2,1): the field is $\mathbb{F}_{2}$ and every acceptance set is a singleton, so each constraint is a linear equation over $\mathbb{F}_{2}$.
We will relate this problem to max-Ekk-LIN-$q$ for some integer $k$ which has been studied more extensively in the literature.
Definition 3 (max-Ekk-LIN-$q$).
Let $\mathbb{F}_{q}$ be the finite field of order $q$ and let $k\geq 1$ be an integer. An instance of max-Ekk-LIN-$q$ consists of $n$ variables $x_{1},\dots,x_{n}\in\mathbb{F}_{q}$ and $m$ linear equations
$\sum_{j\in S_{i}}c_{i,j}x_{j}$ $=b_{i},\qquad i=1,\dots,m,$ (1)
where for each $i$, $S_{i}\subseteq\{1,\dots,n\}$ satisfies $|S_{i}|=k$, $c_{i,j}\in\mathbb{F}_{q}\setminus\{0\}$ for all $j\in S_{i}$, and $b_{i}\in\mathbb{F}_{q}$. The goal is to find an assignment $\boldsymbol{x}\in\mathbb{F}_{q}^{n}$ that satisfies the maximum number of equations.
To establish worst-case limits on the approximation for max-LINSAT, we rely on a foundational result of Håstad [21, Theorem 5.9], which proves strong inapproximability for max-E33-LIN-$\Gamma$ over finite Abelian groups. Håstad’s result builds on the celebrated PCP theorem [4, 3], which characterizes $\mathsf{NP}$ in terms of probabilistically checkable proofs and underlies most modern inapproximability results. This result provides the base case from which we extend hardness to max-LINSAT instances with uniform acceptance sets of arbitrary size.
Theorem 4 (Inapproximability of max-E33-LIN-$\Gamma$).
For every finite Abelian group $\Gamma$ and every $\varepsilon>0$, it is $\mathsf{NP}$-hard to approximate max-E33-LIN-$\Gamma$ within a factor $|\Gamma|-\varepsilon$. Equivalently, max-E33-LIN-$\Gamma$ is non-approximable beyond the random assignment threshold: it is $\mathsf{NP}$-hard to distinguish, given an instance of max-E33-LIN-$\Gamma$, between
  • (Y)
    instances with $\mathrm{OPT}\geq(1-\varepsilon)\,m$ and
  • (N)
    instances with $\mathrm{OPT}\leq(1/|\Gamma|+\varepsilon)\,m$.
In particular, setting $\Gamma=(\mathbb{F}_{q},+)$ for a finite field $\mathbb{F}_{q}$ yields the $(1-\varepsilon,\;1/q+\varepsilon)$-hardness of max-E33-LIN-$q$. Since each coefficient $a_{i}$ is invertible in $\mathbb{F}_{q}$, the substitution $y_{i}=a_{i}x_{i}$ reduces any instance with coefficients to the coefficient-free form (i.e., all coefficients equal to one) covered by the theorem.
The inapproximability guarantee in Theorem 4 is tight. Indeed, for max-E33-LIN-$\Gamma$ over a finite Abelian group $\Gamma$, a random assignment satisfies each equation with probability exactly $1/|\Gamma|$. Therefore, any polynomial-time algorithm can trivially achieve an approximation ratio of $1/|\Gamma|$. Håstad’s theorem shows that it is $\mathsf{NP}$-hard to exceed this baseline by more than an arbitrarily small constant $\varepsilon$, so the result matches the random-assignment threshold up to $\varepsilon$. No stronger worst-case hardness can hold under standard complexity assumptions, making the theorem optimal in this sense.
The inapproximability of max-LINSAT(q,r)(q,r) can also be situated within the broader theory of constraint satisfaction problems. Under the unique games conjecture (UGC) [27], Ref. [36] showed that for every constraint satisfaction problem (CSP), the optimal polynomial-time approximation ratio equals the integrality gap of a canonical semi-definite programming (SDP) relaxation; moreover, this ratio is achieved by an appropriate rounding of the SDP solution. More concretely, Ref. [5] has proven that any predicate whose satisfying assignments support a balanced pairwise-independent distribution is approximation resistant under the UGC. The max-LINSAT(q,r)(q,r) predicate for arity $k=3$, i.e., checking whether $x_{i_{1}}+x_{i_{2}}+x_{i_{3}}$ falls in an $r$-element subset $S\subseteq\mathbb{F}_{q}$, satisfies this criterion (as does the analogous predicate for any fixed $k\geq 3$), so UGC-hardness of beating $r/q$ follows immediately from their result. Our Theorem 5 establishes the same conclusion under the weaker assumption $\mathsf{P}\neq\mathsf{NP}$, via a direct reduction from Håstad’s theorem [21]. This is notable because the UGC remains open. In a different direction, Ref. [16] proved unconditional ($\mathsf{P}\neq\mathsf{NP}$) approximation resistance for predicates whose satisfying assignments contain a pairwise-independent subgroup. For prime $q$, however, the only additive subgroups of $\mathbb{F}_{q}$ are $\{0\}$ and $\mathbb{F}_{q}$ itself, so acceptance sets of size $1\leq r\leq q-1$ are never subgroups, and this criterion does not apply to our setting. For prime power $q$, nontrivial additive subgroups do exist, but the criterion still does not cover generic acceptance sets of arbitrary size $r$. Ref. [18] extended Håstad’s max-E33-LIN hardness from Abelian to all finite groups, and Ref. [13] recently transferred these inapproximability results to the setting of promise constraint satisfaction problems. These results may prove useful for establishing inapproximability bounds for potential DQI variants over more general finite groups.
3 Results
Theorem 5 (Inapproximability of max-LINSAT(q,r)(q,r)).
For every finite field $\mathbb{F}_{q}$, every integer $1\leq r\leq q-1$, and every $\varepsilon>0$, it is $\mathsf{NP}$-hard to distinguish, given an instance of max-LINSAT(q,r)(q,r), between
  • (Y)
    instances with $\mathrm{OPT}\geq(1-\varepsilon)\,m$ and
  • (N)
    instances with $\mathrm{OPT}\leq(r/q+\varepsilon)\,m$.
This means it is $\mathsf{NP}$-hard to approximate within a factor strictly better than $r/q+\varepsilon$ for every $\varepsilon>0$.
Proof.
The case $r=1$ follows directly from Theorem 4, since every instance of max-E33-LIN-$q$ (with coefficients) is syntactically an instance of max-LINSAT(q,1)(q,1): each equation $\sum_{j\in S_{i}}c_{i,j}x_{j}=b_{i}$ defines a constraint with $F_{i}={b_{i}}$.
For general $r\geq 2$, we give a deterministic polynomial-time reduction from max-LINSAT(q,1)(q,1) to max-LINSAT(q,r)(q,r). Let $\phi$ be an instance of max-LINSAT(q,1)(q,1) with $m$ constraints $L_{i}(\boldsymbol{x})=b_{i}$ for $i=1,\dots,m$, where each $L_{i}$ is a linear form over $\mathbb{F}_{q}$. We construct an instance $\phi^{\prime}$ of max-LINSAT(q,r)(q,r) on the same set of $n$ variables as follows. For each constraint $i$, we enumerate all $r$-element subsets of $\mathbb{F}_{q}$ that contain $b_{i}$. There are exactly $\binom{q-1}{r-1}$ such subsets, since we choose the remaining r1r-1 elements from $\mathbb{F}_{q}\setminus\{b_{i}\}$. For each such subset $S$, we add the constraint $L_{i}(\boldsymbol{x})\in S$ to $\phi^{\prime}$. The resulting instance $\phi^{\prime}$ has $m^{\prime}=m\cdot\binom{q-1}{r-1}$ constraints, each with acceptance set of size exactly $r$. Since $q$ is a fixed constant, $m^{\prime}$ is polynomial in $m$.
Completeness. If $\boldsymbol{x}$ satisfies $L_{i}(\boldsymbol{x})=b_{i}$, then $L_{i}(\boldsymbol{x})\in S$ for every $r$-element subset $S$ containing $b_{i}$, so all $\binom{q-1}{r-1}$ derived constraints from constraint $i$ are satisfied. Therefore, $\mathrm{OPT}(\phi^{\prime})\geq\mathrm{OPT}(\phi)\cdot\binom{q-1}{r-1}$. In the (Y)-case, $\mathrm{OPT}(\phi)\geq(1-\varepsilon)\,m$, so that
$\mathrm{OPT}(\phi^{\prime})\geq(1-\varepsilon)\,m^{\prime}.$ (2)
Soundness. Fix any assignment $\boldsymbol{x}\in\mathbb{F}_{q}^{n}$ and define
$A(\boldsymbol{x})$ $=\{i:L_{i}(\boldsymbol{x})=b_{i}\},$ (3)
$\overline{A}(\boldsymbol{x})$ $=\{i:L_{i}(\boldsymbol{x})\neq b_{i}\}.$ (4)
For each $i\in A(\boldsymbol{x})$, all $\binom{q-1}{r-1}$ derived constraints are satisfied. For each $i\in\overline{A}(\boldsymbol{x})$, let $c_{i}=L_{i}(\boldsymbol{x})\neq b_{i}$. The derived constraint $L_{i}(\boldsymbol{x})\in S$ is satisfied if and only if $c_{i}\in S$. Since $S$ must contain $b_{i}$, the number of $r$-element subsets of $\mathbb{F}_{q}$ containing both $b_{i}$ and $c_{i}$ is $\binom{q-2}{r-2}$ (choosing the remaining r2r-2 elements from $\mathbb{F}_{q}\setminus\{b_{i},c_{i}\}$). Thus, for each $i\in\overline{A}(\boldsymbol{x})$, exactly $\binom{q-2}{r-2}$ out of $\binom{q-1}{r-1}$ derived constraints are satisfied. Consequently, the fraction of derived constraints satisfied per unsatisfied original constraint is
$\frac{\binom{q-2}{r-2}}{\binom{q-1}{r-1}}=\frac{r-1}{q-1}.$ (5)
Writing $\mu\coloneqq|A(\boldsymbol{x})|/m$ for the fraction of original constraints satisfied exactly, the total fraction of constraints in $\phi^{\prime}$ satisfied by $\boldsymbol{x}$ is
$\mu\cdot 1+(1-\mu)\cdot\frac{r-1}{q-1}=\mu\cdot\frac{q-r}{q-1}+\frac{r-1}{q-1}.$ (6)
In the (N)-case, Theorem 4 guarantees $\mu\leq 1/q+\varepsilon$ for every assignment. By substituting this and using the identity
$\frac{1}{q}\cdot\frac{q-r}{q-1}+\frac{r-1}{q-1}=\frac{(q-r)+q(r-1)}{q(q-1)}=\frac{r}{q},$ (7)
we obtain
$\mu\cdot\frac{q-r}{q-1}+\frac{r-1}{q-1}\leq\left(\frac{1}{q}+\varepsilon\right)\frac{q-r}{q-1}+\frac{r-1}{q-1}=\frac{r}{q}+\varepsilon\cdot\frac{q-r}{q-1}.$ (8)
Since $(q-r)/(q-1)\leq 1$, this is at most $r/q+\varepsilon$. Therefore $\mathrm{OPT}(\phi^{\prime})\leq(r/q+\varepsilon)\,m^{\prime}$, completing the soundness analysis.
Remark 6 (Tightness of Theorem 5).
The threshold $r/q$ in Theorem 5 is tight. Uniformly chosen assignments $\boldsymbol{x}\in\mathbb{F}_{q}^{n}$ satisfy each constraint $i$ with probability exactly $|F_{i}|/q=r/q$, giving an expected approximation ratio of $r/q$. A deterministic polynomial-time algorithm achieving a ratio of at least $r/q$ is obtained by the method of conditional expectations.
Remark 7 (max-XORSAT).
Setting $q=2$ and $r=1$ in Theorem 5 recovers the result from Ref. [21] stating that max-XORSAT– equivalently, max-E33-LIN-22 – is $\mathsf{NP}$-hard to approximate within $1/2+\varepsilon$ for every $\varepsilon>0$.
4 Discussion
In this work, we have established tight inapproximability results for a family of optimization problems that feature strongly in the decoded quantum interferometry (DQI) algorithm [25], which has been proposed to offer a speedup over classical approaches. Theorem 5 establishes that no polynomial-time algorithm can approximate the problem max-LINSAT(q,r)(q,r) beyond the random assignment threshold $r/q$ on worst-case instances, assuming $\mathsf{P}\neq\mathsf{NP}$, and likewise no polynomial-time quantum algorithm can do so unless $\mathsf{NP}\subseteq\mathsf{BQP}$. In light of the fact that this kind of problem is central to DQI, we now discuss the implications of this result for evaluating quantum optimization algorithms by resorting to DQI and beyond.
A key observation is that DQI’s quantum advantage claims are not for arbitrary max-LINSAT instances, but for algebraically structured problems, most notably optimal polynomial intersection (OPI), where the constraint matrix has Vandermonde structure arising from classical Reed–Solomon codes [25]. For OPI with parameters $n/q=r/q=1/2$, DQI achieves an approximation ratio of approximately $0.933$ via the semicircle law, while the best known classical algorithm (Prange’s information set decoding [35]) achieves only ${\sim}0.55$ [25]. These insights have recently been extended to algebraic geometry codes in Ref. [19], where a problem called HOPI is studied, which generalizes OPI. Therefore, DQI targets structured instances, not worst-case max-LINSAT.
Our worst-case inapproximability result is entirely consistent with these claims: Theorem 5 applies to arbitrary max-LINSAT(q,r)(q,r) instances, whereas OPI instances possess algebraic structure that DQI exploits via coherent implementation of the Berlekamp–Massey decoding algorithm. The hardness reduction from max-E33-LIN-$q$ produces instances with no such exploitable structure.
In DQI, the optimization problem is mapped to a decoding task for a classical error-correcting code. A central parameter governing the algorithm’s performance is the decoding radius $\ell$, defined as the maximum number of errors that an efficient decoder can correct. If a max-LINSAT instance with $m$ constraints arises from a code with decoding radius $\ell$, then DQI coherently maps assignments within Hamming distance $\ell$ to codewords. The resulting approximation ratio is given by the semicircle law [25],
$\alpha_{\mathrm{DQI}}=\left(\sqrt{\frac{\ell}{m}\!\left(1-\frac{r}{q}\right)}+\sqrt{\frac{r}{q}\!\left(1-\frac{\ell}{m}\right)}\right)^{\!2}$ (9)
if $r/q\leq 1-\ell/m$, and $\alpha_{\mathrm{DQI}}=1$ otherwise. This relationship is illustrated in Fig. 1. The ratio $\ell/m$ therefore quantifies the amount of decodable structure available to the algorithm. When $\ell/m$ is bounded away from zero, the decoder can correct a constant fraction of errors, and DQI achieves approximation ratios strictly above the random-assignment baseline $r/q$. In the limit $\ell/m\to 0$, the semicircle law yields $\alpha_{\mathrm{DQI}}\to r/q$: without sufficient decoding capability, DQI reduces to a random assignment.
Figure 1: Landscape of approximability for $\mathrm{max\text{-}LINSAT}(q,r)$. The x-axis parametrizes the decodable structure $\ell/m$, where $\ell$ is the decoding radius of the underlying error-correcting code and $m$ is the number of constraints. The shaded region, spanning approximation ratios from $r/q$ to 11 across all values of $\ell/m$, is $\mathsf{NP}$-hard to achieve on worst-case instances (Theorem 5). The pink curve depicts DQI’s approximation ratio as given by the semicircle law [25], which surpasses the random-assignment bound $r/q$ whenever $\ell/m>0$ and saturates at $\alpha_{\mathrm{DQI}}=1$ for $\ell/m\geq 1-r/q$. In the limit $\ell/m\to 0$, DQI’s performance degrades to exactly $r/q$, matching the worst-case barrier.
This observation is complemented by Theorem 5, which shows that on worst-case max-LINSAT(q,r)(q,r) instances, no polynomial-time algorithm can exceed the $r/q$ threshold (assuming $\mathrm{P}\neq\mathrm{NP}$). While these two statements operate in different regimes – the semicircle law describes DQI’s performance on code-derived instances as a function of the decoding radius, and Theorem 5 applies to adversarially constructed instances with no exploitable structure – they point to the same conclusion from different directions: $r/q$ is the natural worst-case approximation barrier for max-LINSAT, and surpassing it requires both an algorithm capable of exploiting algebraic structure and an instance that possesses such structure. The codes for which DQI demonstrates an advantage over known classical algorithms, namely Reed–Solomon codes in OPI [25] and algebraic geometry codes in HOPI [19], satisfy both conditions.
It is suggestive to read this connection in coding-theoretic terms. For code families where efficient decoding beyond a vanishing fraction of errors is believed to be computationally hard – such as generic linear codes specified by arbitrary parity-check matrices, for which bounded-distance decoding is $\mathsf{NP}$-complete [7] – one would expect the associated max-LINSAT instances to be similarly hard to approximate beyond $r/q$, since a sufficiently good approximate solution would be expected to reveal information about nearby codewords. Making this intuition precise by exhibiting a formal reduction from approximating code-derived max-LINSAT instances to decoding remains an interesting open question.
Recent work has further clarified that DQI’s advantage, where it exists, depends critically on problem structure: Ref. [31] shows that for MaxCut instances where DQI achieves any nontrivial approximation guarantee, the optimal solution can be found exactly in classical polynomial time. DQI consequently provides no advantage for MaxCut. For unstructured max-$k$-XOR-SAT instances, Ref. [2] proves that DQI is obstructed by the overlap gap property, which is a topological barrier in the near-optimal solution space. Classical approximate message passing outperforms DQI on such instances. From the perspective of computational complexity, Ref. [30] places DQI in $\mathsf{BPP}^{\mathsf{NP}}$, thereby ruling out immediate quantum advantage or “supremacy” arguments of the type as they are used in quantum random sampling [20, 8]. There, it is $\mathsf{P}-$hard to compute individual output probabilities, and an efficient classical sampling algorithm up to a constant error in the total variation distance would lead to a collapse of the polynomial hierarchy to the third level. In contrast, DQI is already placed in a low level of the polynomial hierarchy. The hardness of DQI, if any, arises from locating elements in an exponentially large hidden subset without group structure and not from the hardness of simulating the quantum circuit itself, as here, in contrast to the situation in quantum random sampling, individual probabilities of outcomes can be efficiently computed.
Together with Theorem 5, these results rigorously establish that DQI is not a general-purpose quantum optimizer. Several natural questions remain:
  • Inapproximability for OPI and HOPI: Can inapproximability results be extended to settings studied in OPI [25] and HOPI [19]? The Vandermonde structure of OPI instances places them in a restricted subclass of max-LINSAT for which generic PCP reductions do not directly apply, and it remains open whether this structure makes the problem classically easier or merely differently hard. A positive answer would strengthen DQI’s advantage claim; a negative answer (efficient classical algorithms for OPI) would “dequantize” DQI.
  • Average-case hardness: Our result is worst-case. Establishing average-case inapproximability for natural distributions over max-LINSAT instances as they occur in DQI would provide stronger benchmarks for heuristic quantum algorithms.
  • Quadratic constraints: Ref. [23] extends DQI to quadratic constraints over finite fields (max-QUADSAT). For $q=2$ with singleton acceptance sets, Håstad [22] establishes tight inapproximability at the $2^{-d}$ random-assignment threshold for degree-$d$ equations, giving $1/4$ for quadratic constraints, which often appear in practically relevant optimization problems. Extending this to general $\mathbb{F}_{q}$ and acceptance sets of size larger than one remains open: the proof of Ref. [22] relies on $\mathbb{F}_{2}$-specific structure (multilinearity and Reed–Muller minimum weight characterizations) that does not directly generalize.
  • Heterogeneous acceptance sets: Throughout this work (and also in Ref. [25]) the acceptance sets satisfy $|F_{i}|=r$ uniformly. If the cardinalities $|F_{i}|$ are allowed to vary across constraints, a random assignment satisfies constraint $i$ with probability $|F_{i}|/q$, so the natural baseline is no longer a single reference value but depends on the distribution of the $|F_{i}|$. In particular, instances with small acceptance sets ($|F_{i}|=1$) are the most restrictive locally, yet mixtures of small and large sets may exhibit different global approximation behavior. What is a useful notion of a tight inapproximability threshold in this heterogeneous setting? Should it be governed by the minimum $|F_{i}|$, the average acceptance probability $\frac{1}{m}\sum_{i}|F_{i}|/q$, or a more refined structural parameter? Determining an optimal approximation guarantee as a function of the profile $(|F_{i}|)_{i=1}^{m}$ thus remains open.
Our Theorem 5 provides a complexity-theoretic benchmark for max-LINSAT, establishing $r/q$ as the provable worst-case approximation limit and confirming that any improvement must arise from exploitable algebraic structure. It is our hope that this result contributes to clarifying the landscape of quantum advantage for optimization problems of practical relevance [17, 24, 1, 6].
Acknowledgements
This work has been inspired by a recent Computer Science Stack Exchange post [29], which has motivated us to work out the inapproximability of max-LINSAT in full. The authors thank Elies Gil-Fuster for helpful discussions and feedback on this work. Moreover, the authors acknowledge support by the BMFTR (PraktiQOM, QuSol, HYBRID++, PQ-CCA), the Munich Quantum Valley, Berlin Quantum, the Quantum Flagship (Millenion, PasQuans2), the DFG (CRC 183, SPP 2514), the Clusters of Excellence (MATH+, ML4Q), and the European Research Council (DebuQC).